# references.bib

### v1

@@ Abstract
Cognitive Behavioral Therapy (CBT) is a well-established, evidence-based treatment for Major Depressive Disorder. Unfortunately, there exist significant barriers to individuals accessing CBT, including cost, scarcity of therapists and stigma. This study explores the feasibility of fine-tuning small open weight large language models (LLMs) to deliver CBT for depression. Using 58 sets of synthetic CBT transcripts generated by the Nous Research fine-tune of Llama 3.1 405b, we fine-tuned three models: Mistral 7b v0.3, Qwen 2.5 7b, and Llama 3.1 8b. CBT fidelity was evaluated through a modified Cognitive Therapy Rating Scale (CTRS). All fine-tuned models were compared against each other, as well as their instruct-tuned variants. Simulated patient transcripts were generated for the purpose of evaluating model performance, with the instruct and CBT-tuned models acting as the therapist and DeepSeek-V2.5 acting as the patient. These simulated transcripts were evaluated on a modified CTRS by Gemini 1.5 Pro-002. Our findings demonstrated that the CBT-tuned models significantly outperformed their instruct-tuned counterparts, with an average improvement of 11.33 points (p < 0.001) on total CTRS score. Llama 3.1 8b had the strongest performance (mean CTRS score 67.86 ± 7.24), followed by Qwen 2.5 7b (64.28 ± 9.55) and Mistral 7b v0.3 (64.17 ± 9.79), with these differences between models being statistically significant. The CBT-tuned models were competent in implementing core CBT techniques and providing empathetic responses, however, there were limitations observed in agenda adherence, exploration depth and long-context coherence. This study establishes that CBT specific fine-tuning can effectively encode therapeutic competencies in small LLMs, though significant technical and ethical considerations must be resolved prior to clinical deployment.
Keywords: Cognitive Behavioral Therapy, Large Language Models, Depression Treatment, Fine-tuning, Therapeutic AI, Automated Psychotherapy.
1.   Introduction
@@
Major depressive disorder (MDD) is a common psychiatric condition that affects 20.6% of Americans over their lifetime [Hasin2018EpidemiologyStates]. Individuals with MDD experience significant decrements in social functioning, role-emotional functioning, and cognition, resulting in substantial economic burden [Hasin2018EpidemiologyStates, Greenberg2021The2018]. The yearly economic cost of MDD in 2018 was calculated to be \$326.2 billion, significantly increased from \$236.6 billion in 2010 [Greenberg2021The2018]. Despite, the functional impairment and economic loss associated with MDD, many Americans are not able to access evidence-based treatments [Gonzalez2010DepressionFew].
@@
Cognitive behavioral therapy (CBT) is one of the most effective non-pharmacologic interventions for MDD, demonstrating comparable efficacy to antidepressants [Hundt2013TheLiterature, Cuijpers2013ATreatments, Cuijpers2019EffectivenessMeta-analysis, Beck1970CognitiveTherapy, DeRubeis1999MedicationsComparisons]. CBT is a time limited manualized therapy that focuses on identifying and modifying maladaptive thought patterns and behaviors that contribute to symptoms of depression [Beck1970CognitiveTherapy]. Despite the well-established efficacy of CBT in individuals with MDD, it remains underutilized due to a combination of patient, provider and systemic factors. Notable barriers include, fear of stigma, cost, insufficient number of trained therapists and poor access to mental health care in certain geographic regions [Collins2004GapsCare].
@@
Many of the aforementioned challenges could be addressed through the deployment of a CBT delivery system that is powered by artificial intelligence (AI). AI therapists have the potential to offer personalized, scalable and cost-effective psychotherapy to individuals who might otherwise struggle to access face-to-face CBT [Obradovich2024OpportunitiesPsychiatry]. The rapid advancement of large language models has perhaps, for the first time, created an opportunity to craft AI therapists that can deliver true manualized CBT [Obradovich2024OpportunitiesPsychiatry, Stade2024LargeEvaluation]. LLMs excel at text completion, question answering and language understanding [Lappin2024AssessingModels]. Their ability to produce context aware responses and mimic the style of any given template, positions them well to deliver a manualized therapy like CBT [Stade2024LargeEvaluation, Jiang2024ATherapyb].
@@
There have been several papers published within the past two years that have attempted to use LLMs to either augment CBT or provide CBT directly to patients [Shen2024AreTherapy]. Some have used prompt engineering approaches, where an existing closed source model is provided with a specialized prompt to steer its responses to be informed by CBT principles [Iftikhar2024TherapyCBT, Kian2024CanStudents]. HELPERT is one such example, that utilized GPT-4 with a prompt that was designed to guide users through self-reflection and problem solving [Iftikhar2024TherapyCBT]. Other approaches have targeted only a small aspect of the CBT problem space, such as challenging cognitive distortions or Socratic questioning [Izumi2024ResponseQuestioning, Na2024CBT-LLM:Answering].
@@
Unfortunately, there are significant limitations to the previously discussed approaches. Prompt engineering for example, can be brittle and is limited by the context of the model [Knoth2024AIStrategies]. There are often detrimental effects on model coherence when passing too large of a system prompt [Levy2024SameModels]. Furthermore, given that CBT is a manualized therapy where the dialogue and content can vary based on the phase, it would be necessary to engineer multiple prompts for different sessions and situations [Beck1970CognitiveTherapy]. Even with situation specific prompts, it would be difficult to provide a model with all of the information that is required to deliver CBT. Large frontier models such as Claude 3.5 Sonnet, GPT-4o and Llama 3.1 405b would likely not suffer from a lack of knowledge about CBT and may perform well with only prompt engineering. However, most large frontier models are closed source and require significant compute infrastructure for inference [Hoffmann2022TrainingModels]. To our knowledge, there are no papers that have fine-tuned an open weight LLM to deliver a full course of CBT for depression.
@@
In this paper, we propose a method for fine-tuning three small versatile LLMs, Mistral 7b v0.3, Qwen 2.5 7b and Llama 3.1 8b to deliver CBT for depression [Dubey2024TheModels, QwenTeam2024Qwen2.5:Models, Yang2024Qwen2Report, Jiang2023Mistral7B]. Although such a task would be considerably easier if we utilized larger open-weight LLMs, it would run counter to the goal of our approach, which is to contribute to the development of an AI therapist that can deliver cost-effective, accessible and personalized care to individuals with MDD. The training and inference costs associated with highly parameterized models would also exceed our compute budget.
2.   Methods
@@ 2.1 Overview
This study aims to fine-tune small open-weight large language models on synthetically generated CBT transcripts to perform CBT for depression. We start by creating a diverse set of CBT transcripts using the Nous Research fine-tune of Llama 3.1 405b [Teknium2024HermesReport, Dubey2024TheModels]. We fine-tune Mistral 7b v0.3, Qwen 2.5 7b and Llama 3.1 8b on the synthetic transcripts [Dubey2024TheModels, QwenTeam2024Qwen2.5:Models, Yang2024Qwen2Report, Jiang2023Mistral7B]. We then run simulated patient interactions between our fine-tuned models, who act as the therapist, and DeepSeek-V2.5, which acts as the patient [I.2024DeepSeek-V2:Model]. The transcripts that are generated through these simulated therapy sessions are then evaluated on a modified Cognitive Therapy Rating Scale using Google’s Gemini 1.5 Pro-002 [Goldberg2020TheScale, Affrunti2019TheClinicians, Team2024GeminiContext].
@@ 2.2 Synthetic Data Generation
58 sets of CBT transcripts were generated using the Nous Research fine-tune of Llama 3.1 405b [Teknium2024HermesReport, Dubey2024TheModels]. Each set consisted of 20 sessions and was meant to represent a therapy course for one unique patient suffering from MDD. Patient profiles were generated through random selection of characteristics such as age, gender, ethnicity, education level, occupation, symptom severity, engagement level, life events, family background, hobbies, social support systems, personality traits, and coping mechanisms.
@@
Another important aspect of generating synthetic data was to employ the appropriate core CBT techniques across different phases of the therapy [Kazantzis2018TheMeta-Analyses, Fenn2013TheTherapy]. To this end, we divided up the sessions within a course of CBT into four distinct phases: assessment (sessions 1-3), initial (sessions 4-7), middle (sessions 8-17), and termination (sessions 18-20). Each phase was associated with its own goals and CBT techniques. The assessment phase focused on gathering detailed information about the patient’s depression symptoms, life circumstances, and treatment history while beginning to establish the therapeutic relationship. The initial phase aimed to introduce the cognitive conceptualization framework and begin socializing the patient to CBT principles and techniques. The middle phase built upon these principles with deeper exploration and modification of cognitive distortions and maladaptive beliefs. Thought records, behavioral experiments and cognitive restructuring were the main foci of this phase. The termination phase involved consolidating gains from therapy, planning for future challenges and exploring relapse prevention. Our division of CBT phases and associated techniques was informed by the CBT literature [Kazantzis2018TheMeta-Analyses, Fenn2013TheTherapy, McGinn2000CognitiveStatus, Hundt2013TheLiterature, Fennell1987CognitiveChange, Carter2015PredictorsProcess].
Each session within a set was generated in four parts. The first part consisted of a brief check-in with the patient and setting an agenda for the session. The second and third parts focused on discussing previous homework, exploring new techniques relevant to the session phase or building upon techniques introduced in previous sessions. A session concluded with a review of key points, a summary provided to the patient, and homework assigned for the next meeting [Kazantzis2018TheMeta-Analyses].
@@ 2.3 Model Selection and Fine-tuning
The three open-weight models that were selected for fine-tuning were Mistral 7b v0.3, Qwen 2.5 7b, and Llama 3.1 8b. These models were chosen for their robust performance compared to other models of their size, long context windows and wide community support for fine-tuning [Dubey2024TheModels, QwenTeam2024Qwen2.5:Models, Yang2024Qwen2Report, Jiang2023Mistral7B]. We chose to use models in the 7-8 billion parameter range to balance inference costs with model capabilities. We postulated that models smaller than 7 billion parameters would not have the capacity to be fine-tuned for a complex task like CBT, and models in the 20-70 billion parameter range would be outside of our compute budget.
Due to computational constraints, we opted for the Quantized Low-Ranked Adaptation (QLoRA) fine-tuning technique using Unsloth [Dettmers2023QLoRA:LLMs, unslothai2024Unsloth]. We chose the Unsloth library for its built-in memory optimizations which allowed us to fine-tune each model on a single NVIDIA A40 graphics card [unslothai2024Unsloth]. The 4-bit quantized base versions of the models were fine-tuned for one epoch over the training set. To enhance inference fidelity, the fine-tuned adapters were merged with 16-bit versions of the respective base models.
@@ 2.4 Simulated CBT Sessions
The fine-tuned and instruct-tuned versions of Mistral 7b v0.3, Qwen 2.5 7b and Llama 3.1 8b were employed as AI therapists in simulated CBT sessions, with DeepSeek-V2.5 acting as the patient [Dubey2024TheModels, QwenTeam2024Qwen2.5:Models, Yang2024Qwen2Report, Jiang2023Mistral7B, I.2024DeepSeek-V2:Model]. We utilized the instruct-tuned versions of the base models as a point of comparison. For both the CBT fine-tuned and instruct-tuned models, inference was performed with llama.cpp using 8-bit quantized GGUFs [ggerganov2024Llama.cpp].
@@ 2.5 Evaluation Framework
The transcripts from the simulated CBT sessions were graded through an automated evaluation pipeline. Specifically, Gemini 1.5 Pro-002 was used to evaluate the transcripts according to a modified Cognitive Therapy Rating Scale (CTRS). The CTRS is the gold standard for assessing CBT fidelity and involves rating a therapy transcript or recording on the following 11 categories: Agenda, Feedback, Understanding, Interpersonal Effectiveness, Collaboration, Pacing and Efficient Use of Time, Guided Discovery, Focusing on Key Cognitions or Behaviors, Strategy for Change, Application of Cognitive-Behavioral Techniques and Homework [Goldberg2020TheScale, Affrunti2019TheClinicians].
Although the CTRS has only been validated with human therapists as raters, there is a precedent to deploying it with AI systems [Flemotomos2021AutomatedRepresentations]. A 2021 paper used a BERT-based model for automated scoring on the CTRS with encouraging results [Flemotomos2021AutomatedRepresentations]. Our decision to utilize AI raters rather than human raters was informed by multiple factors including rater consistency, scalability and the advancing capabilities of frontier models. Gemini-1.5-Pro-002 was chosen for its excellent recall on long context tasks, state of the art reasoning abilities and free availability through Google’s Gemini API [Team2024GeminiContext]. Furthermore, frontier LLMs such as Gemini-1.5-Pro-002 have been trained on trillions of tokens, resulting in a robust knowledge base in a variety of fields, including CBT [Ke2024ExploringReview].
To increase the resolution of the ratings, the CTRS scale was expanded from a 0-6 Likert scale to a 1-10 scale. Furthermore, given the issues LLMs have experienced when processing numbers and arithmetic, we opted to convert the numeric Likert scale to word-based ratings that ranged from ‘Severely Deficient’ to ‘Outstanding’ [Feng2024HowLLMs]. We acknowledge that the aforementioned modifications to the CTRS will affect its validity in unpredictable ways. However, for the purposes of this pilot study, with minimal human input, we felt that it was appropriate to use the CTRS as a scaffold for our automated evaluation process. For each transcript, Gemini-1.5-Pro-002 was asked to provide both the word-based ratings for each CTRS category as well as justifications for that rating with specific examples. In cases where category evaluation failed after multiple retries, the system employed a conservative approach, assigning the lowest possible score and documenting the failure in the justification. The word-based ratings were then converted back to an integer between 1 and 10 for the statistical analysis.
@@ 2.6 Analysis
Our analysis employed a mixed-effects model, to account for both fixed effects (model type, variant, session number) and random effects (patient-specific variations) [Oberg2007LinearModels]. This approach was chosen to address the nested nature of the data, where multiple sessions were nested within patients, and to account for the non-independence of observations within patient groups. The primary analysis used a linear mixed-effects model with the total CTRS score as the dependent variable. The model was specified as follows:
[…]
In the formula above, session_centered represents the z-standardized session number, model and variant are categorical predictors, and (1 | patient_id) denotes random intercepts for each patient. Fixed effects included the main effects of model type, variant, and session number, without interaction terms. The model was fitted with a maximum likelihood estimate using the LBFGS optimizer [Liu1989OnOptimization]. When convergence issues were encountered, the Powell optimizer was used. [Steihaug2013GlobalFunction].
We implemented the Bonferroni correction method for multiple comparisons of fixed effects [Napierala2012WhatCorrection]. We validated our model by assessing residual normality through Q-Q plots and homoscedasticity using residual versus fitted value plots [Schutzenmeister2012CheckingPlots]. Model fit was assessed using both Akaike Information Criterion (AIC) and Bayesian Information Criterion (BIC) [Kuha2004AICPerformance]. The explained variance was calculated as the ratio between fitted values variance and total score variance.
@@
3.   Results
3.1 Fine-tuning Data
As described in the methods, 58 sets of synthetic CBT transcripts were created using the Nous Research fine-tune of Llama 3.1 405b. Each set of transcripts consisted of 20 sessions, comprising a full course of CBT for depression for a unique patient.

### v2

# Fine-Tuning Open-Weight Language Models to Deliver Cognitive Behavioral Therapy for Depression: A Feasibility Study

@@ Abstract
Cognitive Behavioral Therapy (CBT) is a well-established, evidence-based treatment for Major Depressive Disorder. Unfortunately, there exist significant barriers to individuals accessing CBT, including cost, scarcity of therapists and stigma. This study explores the feasibility of fine-tuning small open weight large language models (LLMs) to deliver CBT for depression. Using synthetic CBT transcripts generated by the Nous Research fine-tune of Llama 3.1 405b, we fine-tuned three models: Mistral 7b v0.3, Qwen 2.5 7b, and Llama 3.1 8b. CBT fidelity was evaluated through a modified Cognitive Therapy Rating Scale (CTRS). All fine-tuned models were compared against each other, as well as their instruct-tuned variants. Simulated patient transcripts were generated for the purpose of evaluating model performance, with the instruct and CBT-tuned models acting as the therapist and DeepSeek-V2.5 acting as the patient. These simulated transcripts were evaluated on a modified CTRS by Gemini 1.5 Pro-002. Our findings demonstrated that the CBT-tuned models significantly outperformed their instruct-tuned counterparts, with an average improvement of 11.33 points (p < 0.001) on total CTRS score. Llama 3.1 8b had the strongest performance (mean CTRS score 67.86 ± 7.24), followed by Qwen 2.5 7b (64.28 ± 9.55) and Mistral 7b v0.3 (64.17 ± 9.79), with these differences between models being statistically significant. The CBT-tuned models were competent in implementing core CBT techniques and providing empathetic responses, however, there were limitations observed in agenda adherence, exploration depth and long-context coherence. This study establishes that CBT specific fine-tuning can effectively encode therapeutic competencies in small LLMs, though significant technical and ethical considerations must be resolved prior to clinical deployment.
1 Introduction
@@
Major depressive disorder (MDD) is a common psychiatric condition that affects 20.6% of Americans over their lifetime [1]. Individuals with MDD experience significant decrements in social functioning, role-emotional functioning, and cognition, resulting in substantial economic burden [1, 2]. The yearly economic cost of MDD in 2018 was calculated to be \$326.2 billion, significantly increased from \$236.6 billion in 2010 [2]. Despite, the functional impairment and economic loss associated with MDD, many Americans are not able to access evidence-based treatments [3].
@@
Cognitive behavioral therapy (CBT) is one of the most effective non-pharmacologic interventions for MDD, demonstrating comparable efficacy to antidepressants [4, 5, 6, 7, 8]. CBT is a time limited manualized therapy that focuses on identifying and modifying maladaptive thought patterns and behaviors that contribute to symptoms of depression [7]. Despite the well-established efficacy of CBT in individuals with MDD, it remains underutilized due to a combination of patient, provider and systemic factors. Notable barriers include, fear of stigma, cost, insufficient number of trained therapists and poor access to mental health care in certain geographic regions [9].
@@
Many of the aforementioned challenges could be addressed through the deployment of a CBT delivery system that is powered by artificial intelligence (AI). AI therapists have the potential to offer personalized, scalable and cost-effective psychotherapy to individuals who might otherwise struggle to access face-to-face CBT [10]. The rapid advancement of large language models has perhaps, for the first time, created an opportunity to craft AI therapists that can deliver true manualized CBT [10, 11]. LLMs excel at text completion, question answering and language understanding [12]. Their ability to produce context aware responses and mimic the style of any given template, positions them well to deliver a manualized therapy like CBT [11, 13].
@@
There have been several papers published within the past two years that have attempted to use LLMs to either augment CBT or provide CBT directly to patients [14]. Some have used prompt engineering approaches, where an existing closed source model is provided with a specialized prompt to steer its responses to be informed by CBT principles [15, 16]. HELPERT is one such example, that utilized GPT-4 with a prompt that was designed to guide users through self-reflection and problem solving [15]. Other approaches have targeted only a small aspect of the CBT problem space, such as challenging cognitive distortions or Socratic questioning [17, 18].
@@
Unfortunately, there are significant limitations to the previously discussed approaches. Prompt engineering for example, can be brittle and is limited by the context of the model [19]. There are often detrimental effects on model coherence when passing too large of a system prompt [20]. Furthermore, given that CBT is a manualized therapy where the dialogue and content can vary based on the phase, it would be necessary to engineer multiple prompts for different sessions and situations [7]. Even with situation specific prompts, it would be difficult to provide a model with all of the information that is required to deliver CBT. Large frontier models such as Claude 3.5 Sonnet, GPT-4o and Llama 3.1 405b would likely not suffer from a lack of knowledge about CBT and may perform well with only prompt engineering. However, most large frontier models are closed source and require significant compute infrastructure for inference [21]. To our knowledge, there are no papers that have fine-tuned an open weight LLM to deliver a full course of CBT for depression.
@@
In this paper, we propose a method for fine-tuning three small versatile LLMs, Mistral 7b v0.3, Qwen 2.5 7b and Llama 3.1 8b to deliver CBT for depression [22, 23, 24, 25]. Although such a task would be considerably easier if we utilized larger open-weight LLMs, it would run counter to the goal of our approach, which is to contribute to the development of an AI therapist that can deliver cost-effective, accessible and personalized care to individuals with MDD. The training and inference costs associated with highly parameterized models would also exceed our compute budget.
2 Materials and methods
@@ 2.1 Overview
This study aims to fine-tune small open-weight large language models on synthetically generated CBT transcripts to perform CBT for depression. We start by creating a diverse set of CBT transcripts using the Nous Research fine-tune of Llama 3.1 405b [26, 22]. We fine-tune Mistral 7b v0.3, Qwen 2.5 7b and Llama 3.1 8b on the synthetic transcripts [22, 23, 24, 25]. We then run simulated patient interactions between our fine-tuned models, who act as the therapist, and DeepSeek-V2.5, which acts as the patient [27]. The transcripts that are generated through these simulated therapy sessions are then evaluated on a modified Cognitive Therapy Rating Scale using Google’s Gemini 1.5 Pro-002 [28, 29, 30].
@@ 2.2 Synthetic data generation
A total of 58 sets of CBT transcripts were generated using the Nous Research fine-tune of Llama 3.1 405b [26, 22], with this number representing the maximum feasible quantity given our compute budget. Each set consisted of 20 sessions and was meant to represent a therapy course for one unique patient suffering from MDD. Patient profiles were generated through random selection of characteristics such as age, gender, ethnicity, education level, occupation, symptom severity, engagement level, life events, family background, hobbies, social support systems, personality traits, and coping mechanisms.
@@
Another important aspect of generating synthetic data was to employ the appropriate core CBT techniques across different phases of the therapy [31, 32]. To this end, we divided up the sessions within a course of CBT into four distinct phases: assessment (sessions 1-3), initial (sessions 4-7), middle (sessions 8-17), and termination (sessions 18-20). Each phase was associated with its own goals and CBT techniques. The assessment phase focused on gathering detailed information about the patient’s depression symptoms, life circumstances, and treatment history while beginning to establish the therapeutic relationship. The initial phase aimed to introduce the cognitive conceptualization framework and begin socializing the patient to CBT principles and techniques. The middle phase built upon these principles with deeper exploration and modification of cognitive distortions and maladaptive beliefs. Thought records, behavioral experiments and cognitive restructuring were the main foci of this phase. The termination phase involved consolidating gains from therapy, planning for future challenges and exploring relapse prevention. Our division of CBT phases and associated techniques was informed by the CBT literature [31, 32, 33, 4, 34, 35].
Each session within a set was generated in four parts. The first part consisted of a brief check-in with the patient and setting an agenda for the session. The second and third parts focused on discussing previous homework, exploring new techniques relevant to the session phase or building upon techniques introduced in previous sessions. A session concluded with a review of key points, a summary provided to the patient, and homework assigned for the next meeting [31].
@@ 2.3 Model selection and fine-tuning
The three open-weight models that were selected for fine-tuning were Mistral 7b v0.3, Qwen 2.5 7b, and Llama 3.1 8b. These models were chosen for their robust performance compared to other models of their size, long context windows and wide community support for fine-tuning [22, 23, 24, 25]. We chose to use models in the 7-8 billion parameter range to balance inference costs with model capabilities. We postulated that models smaller than 7 billion parameters would not have the capacity to be fine-tuned for a complex task like CBT, and models in the 20-70 billion parameter range would be outside of our compute budget.
Due to computational constraints, we opted for the Quantized Low-Ranked Adaptation (QLoRA) fine-tuning technique using Unsloth [36, 37]. We chose the Unsloth library for its built-in memory optimizations which allowed us to fine-tune each model on a single NVIDIA A40 graphics card [37]. The 4-bit quantized base versions of the models were fine-tuned for one epoch over the training set. To enhance inference fidelity, the fine-tuned adapters were merged with 16-bit versions of the respective base models.
@@ 2.4 Simulated CBT sessions
The fine-tuned and instruct-tuned versions of Mistral 7b v0.3, Qwen 2.5 7b and Llama 3.1 8b were employed as AI therapists in simulated CBT sessions, with DeepSeek-V2.5 acting as the patient [22, 23, 24, 25, 27]. We utilized the instruct-tuned versions of the base models as a point of comparison. For both the CBT fine-tuned and instruct-tuned models, inference was performed with llama.cpp using 8-bit quantized GGUFs [38].
@@ 2.5 Evaluation framework
The transcripts from the simulated CBT sessions were graded through an automated evaluation pipeline. Specifically, Gemini 1.5 Pro-002 was used to evaluate the transcripts according to a modified Cognitive Therapy Rating Scale (CTRS). The CTRS is the gold standard for assessing CBT fidelity and involves rating a therapy transcript or recording on the following 11 categories: Agenda, Feedback, Understanding, Interpersonal Effectiveness, Collaboration, Pacing and Efficient Use of Time, Guided Discovery, Focusing on Key Cognitions or Behaviors, Strategy for Change, Application of Cognitive-Behavioral Techniques and Homework [28, 29].
Although the CTRS has only been validated with human therapists as raters, there is a precedent to deploying it with AI systems [39]. A 2021 paper used a BERT-based model for automated scoring on the CTRS with encouraging results [39]. Our decision to utilize AI raters rather than human raters was informed by multiple factors including rater consistency, scalability and the advancing capabilities of frontier models. Gemini-1.5-Pro-002 was chosen for its excellent recall on long context tasks, state of the art reasoning abilities and free availability through Google’s Gemini API [30]. Furthermore, frontier LLMs such as Gemini-1.5-Pro-002 have been trained on trillions of tokens, resulting in a robust knowledge base in a variety of fields, including CBT [40].
To increase the resolution of the ratings, the CTRS scale was expanded from a 0-6 Likert scale to a 1-10 scale. Furthermore, given the issues LLMs have experienced when processing numbers and arithmetic, we opted to convert the numeric Likert scale to word-based ratings that ranged from ‘Severely Deficient’ to ‘Outstanding’ [41]. We acknowledge that the aforementioned modifications to the CTRS will affect its validity in unpredictable ways. However, for the purposes of this pilot study, with minimal human input, we felt that it was appropriate to use the CTRS as a scaffold for our automated evaluation process. For each transcript, Gemini-1.5-Pro-002 was asked to provide both the word-based ratings for each CTRS category as well as justifications for that rating with specific examples. In cases where category evaluation failed after multiple retries, the system employed a conservative approach, assigning the lowest possible score and documenting the failure in the justification. The word-based ratings were then converted back to an integer between 1 and 10 for the statistical analysis.
@@ 2.6 Analysis
Our analysis employed a mixed-effects model, to account for both fixed effects (model type, variant, session number) and random effects (patient-specific variations) [42]. This approach was chosen to address the nested nature of the data, where multiple sessions were nested within patients, and to account for the non-independence of observations within patient groups. The primary analysis used a linear mixed-effects model with the total CTRS score as the dependent variable. The model was specified as follows:
$\text{total\_score}\sim\mathcal{C}(\text{model})+\mathcal{C}(\text{variant})+\text{session\_centered}+(1\mid\text{patient\_id})$ (1)
In the formula above, session_centered represents the z-standardized session number, model and variant are categorical predictors, and (1 | patient_id) denotes random intercepts for each patient. Fixed effects included the main effects of model type, variant, and session number, without interaction terms. The model was fitted with a maximum likelihood estimate using the LBFGS optimizer [43]. When convergence issues were encountered, the Powell optimizer was used [44].
We implemented the Bonferroni correction method for multiple comparisons of fixed effects [45]. We validated our model by assessing residual normality through Q-Q plots and homoscedasticity using residual versus fitted value plots [46]. Model fit was assessed using both Akaike Information Criterion (AIC) and Bayesian Information Criterion (BIC) [47]. The explained variance was calculated as the ratio between fitted values variance and total score variance.
@@
3 Results